# The Great Data Standoff: Researchers vs. Platforms Under the Digital Services Act

### v1

@@ 4. DATA AVAILABILITY AS A PLATFORM GOVERNANCE NARRATIVE
Table 1. Availability of TikTok data for studying systemic risks in Romanian election interference incidents, classified by data category and access mechanism. Each row maps individual data points to their corresponding documentation sources and indicates their accessibility through the Research API and data donations.
@@
We then determine whether each data point is available through the Research API or data donations. The latter was obtained by the authors from their own TikTok research accounts. The API provides qualifying research access to certain public data, while data donations (Van Driel et al., 2021; Zannettou et al., 2023) provide individual users with their personal data, which can then be donated to researchers. While other approaches, such as web scraping, are used for research purposes, we focus on these two mechanisms that comply with the terms of service of the platform and so would render a data access request through Article 40 unnecessary.
Figure 1. Availability of TikTok data for studying systemic risks in Romanian election interference incidents under existing access mechanisms, classified by data type and access method. Nested circles represent data categories with individual data points inside. Colours indicate accessibility, ranging from green (fully available via the Research API or data donations) to purple (entirely inaccessible to researchers), with intermediate shades representing partial access. The analysis highlights critical gaps in access, particularly in user attributes, algorithmic recommendations, and content moderation decisions, which limit researchers’ ability to investigate the systemic risks described in our scenarios.
5. CONCLUSION: REFLECTIONS ON DSA DATA ACCESS FOR THE RESEARCH COMMUNITY

### v2

@@ A Brief Overview of an Election Case Study
We then determine whether each data point is available through the Research API or data donations. The latter was obtained by the authors from their own TikTok research accounts. The API provides qualifying research access to certain public data, while data donations (van2021promises; zannettou2023leveraging) provide individual users with their personal data, which can then be donated to researchers. While other approaches, such as web scraping, are used for research purposes, we focus on these two mechanisms that comply with the terms of service of the platform and so would render a data access request through Article 40 unnecessary.
Figure 1: Availability of TikTok data for studying systemic risks in Romanian election interference incidents under existing access mechanisms, classified by data type and access method. Nested circles represent data categories with individual data points inside. Colours indicate accessibility, ranging from green (fully available via the Research API or data donations) to purple (entirely inaccessible to researchers), with intermediate shades representing partial access. The analysis highlights critical gaps in access, particularly in user attributes, algorithmic recommendations, and content moderation decisions, which limit researchers’ ability to investigate the systemic risks described in our scenarios.
Conclusion: Reflections on DSA Data Access for the Research Community
Paper Checklist
  1. 1.
    For most authors…
    1. (a)
      Would answering this research question advance science without violating social contracts, such as violating privacy norms, perpetuating unfair profiling, exacerbating the socio-economic divide, or implying disrespect to societies or cultures? Yes. This is a conceptual paper that analyses the DSA’s Article 40 data-access mechanism and, via a case study, specifies what kinds of data would be needed to study systemic risks. We do not collect, process, or publish personal data, nor do we scrape platforms or build models that profile individuals or groups.
    2. (b)
      Do your main claims in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes. The abstract and introduction accurately state our contributions.
    3. (c)
      Do you clarify how the proposed methodological approach is appropriate for the claims made? Yes. We clarify that this is a conceptual and methodological paper: we derive data needs from prior literature on systemic risks, map them onto the Romanian election case study, and compare these needs with what TikTok documentation, the Research API, and data donations provide. This approach is appropriate for the claims we make, as our contribution is to outline methodological requirements and gaps, not to produce new empirical findings.
    4. (d)
      Do you clarify what are possible artifacts in the data used, given population-specific distributions? NA. This paper does not use empirical datasets.
    5. (e)
      Did you describe the limitations of your work? Yes. We discuss the limitations of our work, including the interpretational uncertainty of systemic risks under the DSA, the lack of transparency about what data platforms actually collect, and the insufficiency of currently available mechanisms (APIs and data donations) to study election-related systemic risks. We also acknowledge that our case study focuses only on TikTok, not covering the cross-platform nature of systemic risks.
    6. (f)
      Did you discuss any potential negative societal impacts of your work? NA. As a conceptual paper, we do not conduct empirical analysis or release tools or datasets that could have direct negative societal impacts.
    7. (g)
      Did you discuss any potential misuse of your work? NA. The paper does not produce datasets, tools, or empirical models that could be misused. Our contribution is limited to conceptual analysis and a case study framing of data needs under the DSA.
    8. (h)
      Did you describe steps taken to prevent or mitigate potential negative outcomes of the research, such as data and model documentation, data anonymization, responsible release, access control, and the reproducibility of findings? NA.
    9. (i)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? Yes.
  2. 2.
    Additionally, if your study involves hypotheses testing…
    1. (a)
      Did you clearly state the assumptions underlying all theoretical results? NA.
    2. (b)
      Have you provided justifications for all theoretical results? NA.
    3. (c)
      Did you discuss competing hypotheses or theories that might challenge or complement your theoretical results? NA.
    4. (d)
      Have you considered alternative mechanisms or explanations that might account for the same outcomes observed in your study? NA.
    5. (e)
      Did you address potential biases or limitations in your theoretical framework? NA.
    6. (f)
      Have you related your theoretical results to the existing literature in social science? NA.
    7. (g)
      Did you discuss the implications of your theoretical results for policy, practice, or further research in the social science domain? NA.
  3. 3.
    Additionally, if you are including theoretical proofs…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? NA.
    2. (b)
      Did you include complete proofs of all theoretical results? NA.
  4. 4.
    Additionally, if you ran machine learning experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? NA.
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? NA.
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? NA.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? NA.
    5. (e)
      Do you justify how the proposed evaluation is sufficient and appropriate to the claims made? NA.
    6. (f)
      Do you discuss what is “the cost“ of misclassification and fault (in)tolerance? NA.
  5. 5.
    Additionally, if you are using existing assets (e.g., code, data, models) or curating/releasing new assets, without compromising anonymity
    1. (a)
      If your work uses existing assets, did you cite the creators? NA.
    2. (b)
      Did you mention the license of the assets? NA.
    3. (c)
      Did you include any new assets in the supplemental material or as a URL? NA.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? NA.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? NA.
    6. (f)
      If you are curating or releasing new datasets, did you discuss how you intend to make your datasets FAIR? NA.
    7. (g)
      If you are curating or releasing new datasets, did you create a Datasheet for the Dataset? NA.
  6. 6.
    Additionally, if you used crowdsourcing or conducted research with human subjects, without compromising anonymity
    1. (a)
      Did you include the full text of instructions given to participants and screenshots? NA.
    2. (b)
      Did you describe any potential participant risks, with mentions of Institutional Review Board (IRB) approvals? NA.
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? NA.
    4. (d)
      Did you discuss how data is stored, shared, and deidentified? NA.
Appendix A Appendix
@@
Table 1: Availability of TikTok data for studying systemic risks in Romanian election interference incidents, classified by data category and access mechanism. Each row maps individual data points to their corresponding documentation sources and indicates their accessibility through the Research API and data donations.